# Droplet coalescence kinetics: Coalescence mechanisms and thermodynamic non-equilibrium effects under isothermal and non-isothermal conditions

### v1

@@ IV Conclusions and discussion
(i) Coalescence dynamics: In both cases, surface tension drives droplet coalescence, while the resistance arises from an isotropic pressure gradient. However, in the non-isothermal case, the temperature rise due to phase transition latent heat significantly increases the isotropic pressure gradient, though it slightly increases the surface tension gradient near the droplet contact point. This results in a smaller total pressure tensor gradient leading to the deferment of droplet coalescence in the non-isothermal case.
@@
(5) Effect of initial distance: Increasing the initial distance between the droplets delays the coalescence time ($t_{c}$) but does not significantly alter the system’s morphological, non-equilibrium, or entropy production characteristics. The relationships between key times (e.g., $t_{c}$, $t_{\rm Dmax}$, $t_{\dot{S}_{\rm prmax}}$) and $r_{d}$ follow a power-law, with larger exponents in the isothermal case.
This study demonstrates that the dynamics of droplet coalescence are heavily influenced by thermodynamic non-equilibrium effects, particularly in the non-isothermal case. The initial distance between droplets primarily impacts the timing of coalescence, with minimal effect on other characteristics, and the system exhibits similar behavior in its evolution.
Future research will focus on the deformation, fragmentation, cavitation, and jetting processes of droplets under shock wavesYang et al. (2024); Shan et al. (2024); Yang et al. (2025)—phenomena with far-reaching applications. In areas like spray combustion, engine injection, and aerospace, the shape and breakup of droplets are crucial in determining combustion efficiency, thrust, and dynamic performance. From a physical standpoint, this process is marked by increasingly significant thermodynamic non-equilibrium effects, making it an essential domain where kinetic methods are necessary for precise modeling and analysis.
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China (Grant Nos. 11875001 and 12172061), the Hebei Outstanding Youth Science Foundation (Grant No. A2023409003), the Central Guidance on Local Science and Technology Development Fund of Hebei Province (Grant No. 226Z7601G), Science Research Project of Hebei Education Department (Grant No. QN2025184), and the Foundation of the National Key Laboratory of Shock Wave and Detonation Physics (Grant No. JCKYS2023212003).

### v3

@@ IV Conclusions and outlooks
(1) Coalescence dynamics: In both cases, surface tension drives droplet coalescence, while the resistance arises from an isotropic pressure gradient. However, in the non-isothermal case, the temperature rise due to phase transition latent heat significantly increases the isotropic pressure gradient, though it slightly increases the surface tension gradient near the droplet contact point. This results in a smaller total pressure tensor gradient leading to the deferment of droplet coalescence in the non-isothermal case.
@@
(5) Effect of initial distance: Increasing the initial distance between the droplets delays the coalescence time ($t_{c}$) but does not significantly alter the system’s morphological, non-equilibrium, or entropy production characteristics. The relationships between key times (e.g., $t_{c}$, $t_{D\rm max}$, $t_{\dot{S}_{\rm prmax}}$) and $r_{d}$ follow a power-law, with larger exponents in the isothermal case.
This study demonstrates that the dynamics of droplet coalescence are heavily influenced by thermodynamic non-equilibrium effects, particularly in the non-isothermal case. Non-equilibrium measures can provides an earlier and more sensitive means of detecting the dominant processes associated with different mechanisms and their transitions. The initial distance between droplets primarily impacts the timing of coalescence, with minimal effect on other characteristics, and the system exhibits similar behavior in its evolution.
In future work, the following three research directions are proposed to expand and deepen the current study:
(1) Extension beyond the local equilibrium assumption: The current external force are introduced base on the equilibrium distribution function. This assumption requires that the phase separation process proceeds much slower than the relaxation of TNE effects, i.e., $t_{\rm ps}\gg\tau$, where $t_{\rm ps}$ denotes the characteristic time of phase separation. However, as the phase separation process accelerates and the degree of discreteness or nonequilibrium increases, the validity of the local equilibrium assumption becomes questionable. In particular, when the equation of state becomes invalid or unknown, the current modeling approach becomes inadequate. Under such conditions, it becomes necessary to explicitly incorporate intermolecular potential functions to accurately describe molecular interactions and recover the correct thermodynamic behavior Liang et al. (2014, 2018).
(2) Influence of solid walls and shear flows: In microfluidic systems, the presence of solid walls and external shear flows has a significant impact on the droplet coalescence process. Specifically, factors such as the contact angle, wall temperature, and shear strength play critical roles in determining the dynamics of droplet deformation, coalescence, and stabilization. Future investigations should incorporate wetting boundary conditions and shear effects to better capture the complex physical phenomena occurring in confined or structured environments.
(3) Droplet deformation and breakup under shock waves: Future research will focus on the deformation, fragmentation, cavitation, and jetting processes of droplets under shock waves Yang et al. (2024); Shan et al. (2024); Yang et al. (2025)—phenomena with far-reaching applications. In areas like spray combustion, engine injection, and aerospace, the shape and breakup of droplets are crucial in determining combustion efficiency, thrust, and dynamic performance. From a physical standpoint, this process is marked by increasingly significant thermodynamic non-equilibrium effects, making it an essential domain where kinetic methods are necessary for precise modeling and analysis.
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China (Grant Nos. 52278119, 11875001 and 12172061), the Hebei Outstanding Youth Science Foundation (Grant No. A2023409003), the Central Guidance on Local Science and Technology Development Fund of Hebei Province (Grant No. 226Z7601G), Science Foundation of NCIAE (Grant No. ZD-2025-06), Science Research Project of Hebei Education Department (Grant No. QN2025184), and the Foundation of the National Key Laboratory of Shock Wave and Detonation Physics (Grant No. JCKYS2023212003).
Appendix
Table 1 shows definitions, components, and their physical meanings of TNE effects:
Table 1: Definitions, components, and physical interpretations of TNE quantities in DBM.
Quantity Definition Components Physical Meaning
$\bm{\Delta}^{*}_{2}$ $\sum_{i}(f_{i}-f^{eq}_{i})\mathbf{v}^{*}_{i}\mathbf{v}^{*}_{i}$ $xx$, $xy$, $yy$ non-organized momentum flux (NOMF)
$\bm{\Delta}^{*}_{3,1}$ $\frac{1}{2}\sum_{i}(f_{i}-f^{eq}_{i})(\mathbf{v}^{*}_{i}\cdot\mathbf{v}^{*}_{i}+\eta_{i}^{2})\mathbf{v}^{*}_{i}$ $x$, $y$ non-organized energy flux (NOEF)
$\bm{\Delta}^{*}_{3}$ $\sum_{i}(f_{i}-f^{eq}_{i})\mathbf{v}^{*}_{i}\mathbf{v}^{*}_{i}\mathbf{v}^{*}_{i}$ $xxx$, $xxy$, $xyy$, $yyy$ flux of NOMF
$\bm{\Delta}^{*}_{4,2}$ $\frac{1}{2}\sum_{i}(f_{i}-f^{eq}_{i})(\mathbf{v}^{*}_{i}\cdot\mathbf{v}^{*}_{i}+\eta_{i}^{2})\mathbf{v}^{*}_{i}\mathbf{v}^{*}_{i}$ $xx$, $xy$, $yy$ flux of NOEF
$\bar{D}_{2}^{*}$ $\sum{\rho\left|\bm{\Delta}^{*2}_{2}\right|}/{\rho}$ scalar density-weighted average NOMF intensity
$\bar{D}_{3}^{*}$ $\sum{\rho\left|\bm{\Delta}^{*2}_{3}\right|}/{\rho}$ scalar density-weighted average of the flux of NOMF intensity
$\bar{D}_{3,1}^{*}$ $\sum{\rho\left|\bm{\Delta}^{*2}_{3,1}\right|}/{\rho}$ scalar density-weighted average NOEF intensity
$\bar{D}_{4,2}^{*}$ $\sum{\rho\left|\bm{\Delta}^{*2}_{4,2}\right|}/{\rho}$ scalar density-weighted average of the flux of NOEF intensity
$\bar{D}^{*}$ $\sum{\rho\sqrt{\left|{\bm{\Delta}^{*2}_{2}}\right|+\left|{\bm{\Delta}^{*2}_{3,1}}\right|+\left|{\bm{\Delta}^{*2}_{3}}\right|+\left|{\bm{\Delta}^{*2}_{4,2}}\right|}/{\rho}}$ scalar density-weighted average global TNE intensity